OpenStudy (aravindg):

find the eqns of tangents drawn to the curve y^2-2x^3-4y+8=0 from point (1,2)

OpenStudy (aravindg):

@experimentX , @TuringTest , @Mani_Jha , @asnaseer

OpenStudy (mani_jha):

Similar way, Aravind. Find dy/dx at (1,2)-that is the slope of the tangent. Then use the two-point form to get the equation.

OpenStudy (amistre64):

is 1,2 on the curve?

OpenStudy (aravindg):

its a point

OpenStudy (amistre64):

if not, then the question is asking for tangents of the curve that go thru 1,2

OpenStudy (amistre64):

is 1,2 on the cave down or cave up side of the parabola?

OpenStudy (aravindg):

u better check with wolf :P

OpenStudy (amistre64):

err, inside or outside the cave that is

OpenStudy (amistre64):

i aint got to chk it with anything :/

OpenStudy (amistre64):

at x=1; y=3 ... the point 1,2 is under that; and the parabola has a +first coeef; so its on the outside and can intercept at most 2 tangents

OpenStudy (amistre64):

opps, you aint got your exponents lined up do you

OpenStudy (amistre64):

this is a cubic .....

OpenStudy (amistre64):

y^2-2x^3-4y+8=0 2yy' -6x^2-4y' = 0 y' = 6x^2/(2y-4) for the slope of any point in the curve

OpenStudy (aravindg):

thx

OpenStudy (amistre64):

if we shift this and look at the graph from the origin (y+2)^2-2(x+1)^3-4(y+2)+8=0 it looks like we can get a better view of it

OpenStudy (amistre64):

\[y=2\pm\sqrt{2x^3-4}\] \[y' = \frac{3x^2}{\sqrt{2x^3-4}}\] slope of a line from 1,2 to x,y is: \[\frac{y-2}{x-1}= \frac{3x^2}{\sqrt{2x^3-4}}\] sub in for y to get it all in xs \[\frac{\sqrt{2x^3-4}}{x-1}= \frac{3x^2}{\sqrt{2x^3-4}}\] solve for x

OpenStudy (amistre64):

\[2x^3-4 = 3x^3-3x^2\] \[-x^3+3x^2-4 =0\]

OpenStudy (amistre64):

x=-1, 2. but only 2 is in the domain to begin with; so when x=2 we have points of tangency with the line